OpenStudy (anonymous):

a perimeter of a rectangle with width 10 inches and length 13 inches is _______. Be sure to include units.

OpenStudy (anonymous):

the formula for perimeter is P=2w+2l. so to put what you know into that formula, you would have 2*10+2*13=P. Can you solve?

OpenStudy (anonymous):

So it would be 20 + 26 = 46?

OpenStudy (legends):

20+26=46

OpenStudy (anonymous):

correct.